Solve the system: x - 5y = -9
x = y - 1
One of the solutions is :
a) x = 1
b) x = -1
c) y = -2
d) y = 2
e) none of the above

Answer by funmath(2933)

Solve the system: x - 5y = -9
x = y - 1
One of the solutions is :
a) x = 1
b) x = -1
c) y = -2
d) y = 2
e) none of the above
:
Substitute y-1 in for x in the first equation and solve for y:
(y-1)-5y=-9
-4y-1=-9
-4y-1+1=-9+1
-4y=-8
-4y/-4=-8/-4
y=2
Substitute 2 in for y in the second equation and solve for x.
x=2-1
x=1
The solution is (x,y)=(1,2)
You need both an x and y for a solution. so the answer is (a and d) or e.
Happy Calculating!!!